These are the news starting 01 Jan 2015
Click here for the news of the previous years.
|16 Dec 15||Maintenance by T.S.E|
|01 Dec 15||Upload of nov15.dat file into data archive.|
|21 Nov 15||Maintenance and calibration by Raoul Tholl.|
|05 Nov 15||There was an interruption from 04Nov15 10:30 to today 13:00 due to a problem with the serial COM port of the datalogger computer (due to a timing conflict, the port was not released and the half-hour transfers were blogged). This problem has been solved (13:40).|
|01 Nov 15||Uploaded oct15.dat into data archive. Logger memory cleared.|
|28 Oct 15||Maintenance and recalibration.|
|16 Oct 15|| Maintenance by T.S.E.
Remember: The WBGT sensor was shutdown the 17th September; will be reactivated in May 2016. Normal icons have been replaced by "not measured" icons.
|07 Oct 15||A group of students from our Japanese "sister" school at Kijimadaira visited the meteorological station. See pictures here: 1, 2, 3, 4. (all pictures by Frank Schreiner).|
|01 Oct 15||Uploaded sep15.dat into data archive.|
|17 Sep 15|| 1. Maintenance on both UV sensors: replaced
dessicant capsule, checked Peltier temperatures (both were bad). Bad
UVB Peltier temp. solved by cleaning connector (bad UVB_Tin during
last weeks probably does not mean that the actual temp. was not
correct, but the Temp. signal was degraded). UVA_Tin still a bit low
(23°C instead of 25°C), not a big problem..
2. Replaced dirt filter of rain sensor.
3. Made numerous backup images of the different computers.
4. Stopped heat-stress WBGT sensor for this year.
|01 Sep 15||Uploaded aug15.dat into data archive.|
|28 Aug 15||Maintenance and calibration work.|
|14 Aug 15||The WOUDC has changed its method to access the data from the contributing stations. The old file showing how to do this has been changed to reflect the new procedure. You find the link on the home page, or if curious, click here.|
|13 Aug 15||The "DOBSON" page has been updated with two graphs comparing the year 2014 direct sky Dobson measurements of Uccle and Diekirch. Our readings are not too bad, but have a low bias: a multiplier of ~1.04 is suggested to adjust them to the Uccle readings.|
|12 Aug 15||A comparison between the two Microtops #3012 (the older model used by Mike Zimmer) and #5375 (the newer model used by Francis Massen) shows that the old model is consistently higher. Starting 01Aug15, a multiplier of 0.8822 will be used with the #3012 readings. Remember that the DOBSON page always holds the readings of instrument #5375 (magenta color) or the adjusted readings of #3012 (blue color).|
|01 Aug 15||Uploaded jul15.dat into data archive.|
|27 Jul 15||Maintenance and calibration.|
|14 Jul 15||Mounted an A-21ZX small ozone sensor from Ecosensors on the terrasse, to check if this very imprecise handheld instrument could be used as an emergency backup... probably not!|
|03 Jul 15||Uploaded jun15.dat into data archive. Watch remark concerning the new SMP10 solar sensor.|
|30 Jun 15||As we have our first very warm day, it is a good idea to remember what heat stress is and how it is measured. Read my blog comment on this problem.|
|26 Jun 15||Calibration tests on the CO2 sensor.|
|19 Jun 15||The AALCD (Amicale des Anciens du Lycée Classique de Diekirch) visited meteoLCD and our Computarium. I gave a short presentation on meteoLCD (which comes to 20 years of operation this year!) and the Computarium. Raoul Tholl made the outside visit and explained the different sensors.|
|18 Jun 15||I installed a box with an old VALTRONICS CO2 sensor (which works by diffusion) together with an Onset logger which will log the CO2 readings (actually a voltage) and ambient air temperature. The aim is to see if these readings could be useful to replace the API CO2 sensor in event of a complete breakdown of the API. Also installed for intercomparison over several months are two NIPHARGUS small temperature loggers from our Belgian speleological friends; these sensors are in the small blue watertight cylinders. See here and here.|
|16 Jun 15|| 1. Maintenance by T.S.E. Again an nagging problem with
our CO2 calibration.
2. Definitive mounting of the SMP10 pyranometer by technician Jean-Paul Klein and me. Click here for a picture of Jean-Paul drilling through 10mm solid steel.
|05 Jun 15|| 1. The new SMP10 pyranometer has been installed
(definitive fixation and final electrical wiring will be done later)
and is running. For a few weeks, a plot will be made containing both
pyranometer readings; the data of the new SMP10 are in the last
column of the data file (today this column still contains some
garbage at the start of the day, but this will be ok tomorrow). The
comparison plot can be seen as the last plot at the right side of
the screen (right of cell with the words "TRENDS").
When everything is ok, the SMP10 data will be moved to the place of the file containing the CM21 readings, and the plot accordingly. The CM21 will than be moved to the last column of the data file and not be plotted anymore. The CM21 remains installed and running, and will be a backup if something goes wrong with the new pyranometer. Click here for a picture of the new instrument (in front, with the curiously shaped fan housing and the yellow cables; the old CM21 is at the back)
2. This evening we had our first thunderstorm of the year. Our old Boltek sensor (together with one of the latest Lightning2000 software) behaved quite well and showed the storm moving from south-west to north-east (the usual direction), with the southern part untouched by the storm. Click on the small pictures below; the EUCLID situation is 2 hours previous, as Euclid takes two hours to put its lightning pictures on the web. The Blitzortung side is live, but tiny Luxembourg remains invisble behind the lightning symbols.
There remains one glitch at the lightning page: the "today" link is not properly updated, but the "last hour" link is ok. I also must take the time to look at all older movies and to store them into the yearly folders.
|04 Jun 15|| Two links at the live-data page concerning ground ozone
in Germany and EU have been changed. The German link is most useful
if one selects first "Ozon" and "Rheinland-Pfalz", and then the
station of TRIER.
The new EU-EEA map also needs some efforts to become acquainted with... I have my doubt on the validity of the hourly data chart (I think this chart represents the daily average...).
|03 Jun 15|| 1. Maintenance work and recalibration of CO2 sensor with
new bottle of calibration gas. There could be a remaining problem
with the sample pump.
2. Uploaded may15.dat file into data archive.
|27 May 15||Upload of original Microtops II (instrument #5375) raw data into the data archive. The file name is 27MAY15.mt5. These are absolute ***RAW*** and ***ORIGINAL*** data, as stored in the instruments memory (which will be cleared to make room for new measurements).|
|04 May 15||Upload of April 2015 data into archive.|
|01 May 15||Maintenance by Raoul Tholl.|
|10 Apr 15|| 1. The WBGT heat-stress sensor has been reinstalled.
Look at a small
album of the equipment.
2. Maintenance of the the CO2 and O3 sensors and overall cleaning. The LNI zero air generator is down again and has been shut off. This should not have consequences on the measurements, as the zero air is used to purge an optical path of the CO2 sensor which might be influenced by people in the intrumentation room. As there are none...
3. The web-cam has been shut off for an indefinite time, due to a proxy failure, very out-dated Java software and general legal uncertainty.
|01 Apr 15||Uploaded March 2015 data into data archive. Logger memory cleared.|
|20 Mar 15|| 1. Maintenance by T.S.E. The air pump of the CO2 sensor
broke down again at the 15th March. This pump has now definitively
been replaced by a bigger non-standard model. Let's hope that these
annoying and recurrent pump problems are now past history.
2. The partial solar eclipse was a moot event at Diekirch: heavy fog with a totally invisible sun was present until the last half-hour. I had prepared a telescope on the meteoLCD terrace, and also a datalogger to make fast (dt = 10 seconds) measurements of solar irradiance, temperature and humidity. Look here for some pictures and more details.
|01 Mar 15||Uploaded February 2015 data into data archive. Edit of "index.html" file of data archive which was not up to date.|
|22 Feb 15|| I added (if not yet present) a worksheet "dailys" to
all data files of the type "yyyy_only.xls" (e.g.
2003_only.xls). This worksheet holds the daily averages of a subset
of instruments and also the Hellmann numbers. These numbers are used
to classify winters as more or less severe (google "Hellmann
numbers" or "Hellmann Zahlen" for explanation). Hellmann defines
winters as the period from 1 November to 31 March; this is not the
usual WMO definition of Dec-Jan_Feb !
If you want to add daily averages of other instruments, copy the formula used in the first row and copy it down.
|14 Feb 15||Maintenance and calibration work.|
|01 Feb 15||Uploaded January 2015 data into data archive. Logger memory cleared.|
|17 Jan 15||Maintenance and calibration work. The trends section is completely updated to include the 2014 data.|
|03 Jan 15||I started today updating the trends page with data from 2014. This is a major work, requiring a lot of calculations and miscellaneous research, so it will be a "work in progress". The updated parameters are clearly marked to avoid any confusion.|
|01 Jan 15|| 1. Overplot of total ozone column measurements done at Uccle and
meteoLCD now on the Dobson page (here).
This gives a first impression of the similarity of both measurement
2. Next 7 days graphs will be corrupted due to the usual Gnuplot problem. This proble is sel-healing in 7 days:-)))
3. Upload of December 2014 data and full year 2014.zip file into data archive.
4. Update of the "Global Warming Sceptic" page with archiving of the 2013-2014 contents.
5. Excel file 2014_only.xls with all readings and many statistics (not fully finalized, as Dec2014 Findel data are still missing!) and file meteoLCDstats_1998_2014.xls uploaded into data archive.